The following fractions represent just three different numbers. Separate them into three groups of equivalent fractions, by changing each one to its simplest form.
12/60

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

12 / 60 = (1 × 12) / (5 × 12)
=1 / 5

​Thanks
Shraddha Singh
2 years ago
Dear Student,
 
The simplest form of the fraction 12/60 is
12/60 = 12÷12/60÷12 = 1/5
 
Hope it helps!
Thank you ✌️